OpenStudy (anonymous):

what is f(5) if f(1) = 3.2 and f(x + 1) = (f(x))

OpenStudy (isaidavila):

F(1+1)=f(1) F(2)=3.2 F(2+1)=f(2) F(3)=3.2 Do the same till you reach f(5)

OpenStudy (unklerhaukus):

f(x+1) = f(x) f(5) = f(4) = f(3) = ...